# Efficient Bayesian Updates for Deep Learning via Laplace Approximations

### v2

@@ Appendix F Active Learning: Use Case
Algorithm: The main idea from Section 5 is to transform any sequential selection strategy into a batch strategy by using the proposed update method as a fast alternative to retraining. Typically, when using a sequential selection strategy, the naive idea in a batch setting is to select the instances that resulted in the top-$b$ highest scores, where the score is defined by the respective selection strategy. Our method can now update the DNN after each acquired label instead of selecting the top-$b$ instances. Hence, we have a similar scenario as if we would perform single-instance acquisitions but avoid the retraining after each acquired label. After aggregating a batch of $b$ instances, we retrain the DNN.
@@
Algorithm 1 summarizes this idea. Given a standard AL setting with labeled data $\mathcal{L}$ and unlabeled data $\mathcal{U}$, we first calculate the approximate posterior distribution $q(\boldsymbol{\omega}|\mathcal{L})$ via LA. Subsequently, we iteratively select a new instance ${\boldsymbol{x}}$ and acquire its label $y$ based on a given selection strategy $\alpha(\cdot)$ and update the approximate posterior distribution $q(\boldsymbol{\omega}|\{({\boldsymbol{x}},y)\},\mathcal{L})$ using our method. We repeat this for $b$ acquisitions.
@@
6:     Extend new acquisitions $\mathcal{D}^{\oplus}\leftarrow\mathcal{D^{\oplus}}\cup\{(\hat{{\boldsymbol{x}}},\hat{y})\}$
@@
Experiments: For the AL experiments in Section 5, we follow the recent work of Hacohen et al. [16]. We use the CIFAR-10 dataset and initialize the labeled pool $\mathcal{L}$ with $10$ randomly sampled instances. In each AL cycle, we select $b=10$ new instances for labeling until a budget of $B=200$ acquisition is reached. Each AL selection strategy selects from $1{,}000$ randomly sampled instances from the unlabeled pool $\mathcal{U}$ to speed up the selection process. We employ the same architecture and training hyperparameters as described in the experimental setup in Section 4.1. For the sequential selection strategies Margin [3] and Bald [14], we select batches of instances by following Algorithm 1 instead of the top-$b$ selection. In the case of Badge [1], we replace the $k$-means++ algorithm that is typically used. Consequently, we select the instance with the highest gradient norm. The learning curves in Fig. 10 show that the proposed idea improves all investigated selection strategies.
Accuracy
Number of Labels
Refer to captionNumber of Labels
Refer to captionNumber of Labels
Figure 10: Learning curves reporting the accuracy of AL strategies combined with our update.

### v3

@@ Appendix F Active Learning: Updates for Batch Selection and Look-Ahead
Algorithm: The main idea from Sec.˜5 is to transform any sequential selection strategy into a batch strategy by using the proposed update method as a fast alternative to retraining. Typically, when using a sequential selection strategy, the naive idea in a batch setting is to select the instances that resulted in the top-$b$ highest scores, where the score is defined by the respective selection strategy. Our method can now update the DNN after each acquired label instead of selecting the top-$b$ instances. Hence, we have a similar scenario as if we would perform single-instance acquisitions but avoid the retraining after each acquired label. After aggregating a batch of $b$ instances, we retrain the DNN.
@@
Algorithm˜1 summarizes this idea. Given a standard AL setting with labeled data $\mathcal{L}$ and unlabeled data $\mathcal{U}$, we first calculate the approximate posterior distribution $q(\boldsymbol{\omega}|\mathcal{L})$ via LA. Subsequently, we iteratively select a new instance $\boldsymbol{x}$} and acquire its label $y$ based on a given selection strategy $\alpha(\cdot)$ and update the approximate posterior distribution $q(\boldsymbol{\omega}|\{({\boldsymbol{x}},y)\},\mathcal{L})$ according to Eq.˜10. We repeat this procedure for $b$ acquisitions.
@@
[…]
@@
Experiments: For the AL experiments in Sec.˜5, we follow the recent work of [16, 15]. We use all datasets (cf. Tab.˜1) and initialize the labeled pool $\mathcal{L}$ with $b$ randomly sampled instances. In each AL cycle, we select $b$ new instances for labeling and fix the number of AL cycles to $20$, resulting in a total budget of $B=20\cdot b$. The size $b$ is determined based on the dataset’s complexity, ensuring that learning curves from random instance selection converge. Each AL selection strategy selected from 1,0001{,}000 randomly sampled instances from the unlabeled pool $\mathcal{U}$ to speed up the selection process, except for Typiclust, where we increased the number to $10,000$ to avoid errors of $k$-means. We employ the same architecture and training hyperparameters as described in the experimental setup in Sec.˜4.1.
Improved Batch Selection: For the sequential selection strategy Margin [3], we select batches of instances by following Algorithm˜1 instead of the top-$b$ selection. In the case of Badge [1], we replace the $k$-means++ algorithm that is typically used. Consequently, we select the instance with the highest gradient norm. All accuracy improvement curves are shown in Fig.˜9. Associated learning curves reporting the absolute accuracy are shown in Fig.˜10.
Figure 9: Accuracy improvement curves for different strategies and datasets showing the accuracy difference between the respective selection strategy and random instance selection.
Figure 10: Learning curves for different strategies and datasets showing the accuracy.
Approximating Optimal Batch Selection: In Sec.˜5, we employed our update to approximate an optimal look-ahead batch selection strategy. The selection works as follows: We begin with a randomly initialized labeled pool. During selection, we deliberately avoid clustering schemes to simplify batch selection into picking a single instance per cluster. We argue that this enforced diversity can lead to suboptimal selection, especially in the later stages of active learning, where exploitation is more critical. Instead, we randomly sampled 2,0002{,}000 batches, each matching the acquisition size. These batches likely include both diverse and non-diverse sets, allowing for exploration in the beginning and potential exploitation in the end. Each batch is used to update the model with our proposed method as a proxy for retraining. We then evaluated the performance of each updated model on a validation set. Finally, we selected the batch that yielded the highest performance improvement. The remaining accuracy improvement curves and learning curves reporting absolute values are shown in Figs.˜11 and 12, respectively.
Figure 11: Accuracy improvement over random selection of popular selection strategies compared to our upper baseline approximating optimal batch selection.
Figure 12: Learning curves reporting the accuracy of popular selection strategies compared to our upper baseline approximating optimal batch selection.